rec-mode: an Emacs mode for editing recfiles

for version 1.8.90, 19 June 2022

by Jose E. Marchesi

This manual is for rec-mode 1.9.0 for Emacs, which is part of the GNU recutils suite. This rec-mode manual has last been updated 19 June 2022. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2012-2022 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Activation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Feedback . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Navigation mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.1 Record navigation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Field navigation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 Field folding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Field edition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.5 Searches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.6 Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.7 Data integrity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3 Edition modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4 Summary mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

5 Selecting records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

5.1 Buffer from selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 5.2 Selection mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

6 Cross referencing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6.1 Cross-referencing records using foreign keys . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.1.1 Navigating to referred records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 6.1.2 Finding record references. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 6.1.3 Commands for cross-referencing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

6.2 Cross referencing by selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

7 Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

7.1 Finding the recutils . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 7.2 Records appearance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Command and Function Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Key (Character) Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

Variable and Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

Appendix A GNU Free Documentation License . . . . . . . . . . . . . . 15

1 Introduction

rec-mode is a mode for browsing and editing recfiles, which are text files containing data structured in fields and records. It is part of the GNU recutils^{1} suite.

Recfiles are text-based databases which are easy to read and write manually using a text editor. At the same time they feature enough structure so they can be read, edited and processed automatically by programs.

1.1 Installation

rec-mode is implemented in a self-contained elisp file called rec-mode.el. It can be obtained in several ways:

 − From GNU ELPA (https://elpa.gnu.org/packages/rec-mode.html) using the built-in

Emacs package manager. − As part of the source tree cloned from the development git repository of rec-mode.

rec-mode.el can be found in the root directory of the recutils sources tree. − As a single file downloaded form some other location in internet. − It may be already installed as part of a binary package in some distribution.

In the second case you need to tell Emacs where to locate the rec-mode.el file and to load it. Add the following to your .emacs file.

     (add-to-list 'load-path "~/path/to/recmode/")
     (require 'rec-mode)

If rec-mode.el was installed as part of a binary package in a distribution then you usually don’t have to touch the load-path variable. Depending on the specific case you may have to require the package.

1.2 Activation

rec-mode should automatically activate when opening .rec files. To make sure files with extension .rec use rec-mode, add the rec-mode buffers need font-lock to be turned on - this is the default in Emacs^{2}.

If the file doesn’t end in .rec, make the first line of a recfile look like this:

     # -*- mode: rec -*-

which will select rec-mode for this buffer no matter what the file’s name is.

1.3 Feedback

If you find problems with rec-mode, or if you have questions, remarks, or ideas about it, please mail to the recutils mailing list bug-recutils@gnu.org. If you are not a member of the mailing list, your mail will be passed to the list after a moderator has approved it^{3}.

1 http://www.gnu.org/software/recutils 2 If you don’t use font-lock globally, turn it on in the rec buffer with (add-hook 'rec-mode-hook 'turn-on-

font-lock)

3 Please consider subscribing to the mailing list, in order to minimize the work the mailing list moderators have

to do. The subscription can be done online at http://lists.gnu.org/mailman/listinfo/bug-recutils.

2 Navigation mode

When a recfile is visited in Emacs and rec-mode is activated, the contents of the file are examined and parsed in order to determine if it is a valid recfile and, in that case, to extract information like the kind of records stored in the file.

If the file does not contain valid rec data then the buffer is put in fundamental-mode and a description of the syntax error, along its location, is notified in the echo area.

If the file contains valid rec data, the mode sets itself in what is known as “navigation mode”. In this mode the buffer is made read-only and it is narrowed to the first record present in the file. Also, the presentation of the record contents is slightly changed in order to improve the visualization of the data: continuation line marks are replaced by indentation, big fields are folded, etc. The modeline is changed in order to reflect the type of the records being navigated.

At this point the user can navigate through the records and fields contained in the file, and edit the contents of the fields and the structure of the records, by using the commands described in the following subsections.

2.1 Record navigation

The following commands jump to other records in the buffer.

n
M-x rec-cmd-goto-next-rec

Display the next record of the same type in the buffer.

‘C-u N n’ will move next N times.

p
M-x rec-cmd-goto-previous-rec

Display the previous record of the same type in the buffer.

‘C-u N p’ will move backwards N times.

d
M-x rec-cmd-show-descriptor

Display the record descriptor applicable to the current record. If the current record is anonymous, i.e. there is not record descriptor. then this command does nothing.

b
M-x rec-cmd-jump-back

Display the record previously displayed in the buffer.

C-c C-t
M-x rec-find-type

Prompt the user for one of the record types present in the recfile and display the first record of the selected type.

2.2 Field navigation

The following commands iterate through the fields in a record, and to get information about some of the properties of the fields.

TAB
M-x rec-cmd-goto-next-field

Move the cursor to the beginning of the name of the next field in the current record. If the cursor is currently located at the last field of the record then move it to the beginning of the first field.

t
M-x rec-cmd-show-type

Show information about the type of the field under the cursor, if it is defined.

2.3 Field folding

Fields in recfiles can contain data of any size, and sometimes it is difficult to have an overview of the contents of the record. The following commands fold and unfold the value of the field under the cursor.

SPC
M-x rec-cmd-toggle-field-visibility

Toggle the visibility of the field under the cursor. When a field is folded then three dots are displayed in the buffer instead of the value of the field. It is possible to automatically fold any field whose value exceeds a certain limit which can be configured by the user. See Section 7.2 [Records appearance], page 11.

2.4 Field edition

The following commands change the value of the field under the cursor.

e
M-x rec-cmd-edit-field

Edit the value of the field under the cursor. The specific action depends on the type of the field in the corresponding record descriptor: − For date fields a calendar buffer is opened in another window and the focus is moved there. The user can then select a date by moving the cursor there and press RET in order to set that date as the value for the field. Alternatively the user can press t in order to set the field to “now”, or q to cancel the operation. In the later case the value of the field is left untouched.

− For enumerated and bool fields a fast-select buffer is opened in another window, showing a list of labeled options. The labels are single digits and letters. The user can then select ony of the options by pressing the corresponding label, or cancel the operation by pressing RET. In the later case the value of the field is left untouched.

− For any other kind of fields an edition buffer is opened in another window, showing the current contents of the field. The user can then edit the buffer as desired. When she is done, the user can then press C-c C-c in order to set the new value of the field, or just kill the buffer to cancel the operation.

m
M-x rec-cmd-trim-field-value

Trim the value of the field under the cursor, removing any sequence of leading and trailing blank characters.

2.5 Searches

The following commands jump to the first record in the buffer satisfying some criteria.

s q
M-x rec-cmd-jump-to-fast

Display the first record having a field whose value matches a given fixed pattern. This is equivalent of using the command line option -q of recsel. If a prefix argument is specified then the search is case-insensitive.

s s
M-x rec-cmd-jump-to-sex

Display the first record in the buffer satisfying a given selection expression. This is equivalent of using the command line option -e of recsel. If a prefix argument is specified then the search is case-insensitive.

2.6 Statistics

The following commands allow to count records in the current buffer based on some provided criteria.

I
M-x rec-cmd-show-info

Show the number of records in the buffer categorized by type.

#
M-x rec-cmd-count

Count the number of records in the buffer having the same type as the current record. With a numeric prefix N, ask for a selection expression and count the number of records in the buffer satisfying the expression. Note that rec-mode tries to guess a reasonable default for the selection expression, depending on the type of the field and its value. If the user press RET then the provided default selection expression is used.

%
M-x rec-cmd-statistics

If the field under the cursor contains an enumerated value, show the percentages of records in the current record set having fields with each of the possible values of the enumerated type.

2.7 Data integrity

The integrity of the rec data stored in the file can be checked using the following commands.

c
M-x rec-cmd-compile

Compile the buffer with recfix and open a compilation window showing the result of the command. In case some error or warning is reported, the user can jump to the location triggering the error by pressing RET in the compilation window.

3 Edition modes

The navigation mode described in a previous chapter is mainly intended for browsing recdata and doing changes at the record level: editing the contents of a field, adding or removing fields, etc. In order to perform broader changes, such as adding/deleting record descriptors, records or comment blocks, the user must enter into one of the “edition modes”.

There are three edition modes, covering different areas of the recfile: record, record type and buffer. When an edition mode is entered the buffer is set in read/write mode, it is narrowed to the desired area and any embellishment used in navigation mode is removed^{1}. As a general rule, the commands available in navigation mode are also available in the edition mode prefixed with C-c and with a control sequence C-. Thus, n, which in record navigation mode jumps to the next record, becomes C-c C-n in the record edition modes. The exceptions to this are the

bindings c (rec-cmd-compile) which becomes C-c C-k and h (rec-cmd-show-summary) which

becomes C-c M-h.

The following commands are used to enter into one of the available edition modes from the navigation mode.

R
M-x rec-edit-record

Edit the record being navigated.

T
M-x rec-edit-type

Edit the record set being navigated.

B
M-x rec-edit-buffer

Edit the buffer.

After doing modifications in the buffer, the user can go back to navigation mode by using the following command.

C-c C-c
M-x rec-finish-editing

Finish the current edition and return to navigation mode. If a syntactic error was introduced in the edition activity then the error is reported in the echo area and navigation mode is not entered.

1 Exceptuating font-lock

4 Summary mode

Besides the standard viewing modes of buffers, there is also a third option called summary mode. Summary mode is a tabulated list of the record in which each record is a row in the table. The columns in the table are determined by user input, or alternatively, by special keys in the current file.

h
M-x rec-cmd-show-summary

Creates a summary buffer from the current type being viewed in the recfile. The columns are based on user input (prompted), or if the record has either a %summary or %key in its descriptor, those fields are used.

The input for the summary query is a field expression, a comma-separated list of column names. Thus, if your recfile looks like the following:

     Name: Ada Lovelace
     Age: 36
     Name: Peter the Great
     Age: 53
     Name: Matusalem
     Age: 969

Then calling rec-cmd-show-summary and entering Name,Age would create a table of all records with those columns. If your record descriptor contains a %summary like below:

     %summary: Age,Name

then this is used as the field expression instead. Similarly, if no %summary field is defined and a %key field exists, then that field is used.

Calling this command enters record summary mode, which a a major mode derived from tabulated list mode (see Section “Tabulated List mode” in elisp). In this mode, the following bindings are available:

n
M-x rec-summary-cmd-next

Go to the next record in the summary.

Navigation inside the summary buffer maintains the position of the associated recfile buffer, and vice versa. The currently active record in the recfile is shown highlighted in the buffer. As a result, the summary can be thought of as a useful “overview” overview for any recfile buffer.

5 Selecting records

Sometimes it is useful to operate or inspect only a subset of all the records on the file. To this end, rec-mode provides several ways of using the searching facilities of recutils.

Selection can be done using two mechanisms:

• selection expressions, which are expressions that can be applied to a record. These offer a powerful query language for records. See Section “Selection Expressions” in GNU Recutils manual.

• fast string searches, which match records whose field values match a string.

For instance, you can create a new buffer out of the current buffer, so that only the records that match your search are copied. This way you can create new recfiles out of existing ones based on selections.

To help navigate the buffer based on a search, you can create a new navigation mode where the commands for moving between records only move to those matching a search. This is known as selection mode.

5.1 Buffer from selection

Creating new recfiles based on selections can be done by creating a wholly new buffer using the buffer from selection commands in rec-mode.

x s
M-x rec-cmd-new-buffer-from-sex

Creates a new buffer from a selection expression. See Section “Selection Expressions” in GNU Recutils Manual. Prompts the user for a selection expression and then displays a new buffer containing records that match the expression.

x q
M-x rec-cmd-new-buffer-from-fast-string

Creates a new buffer from a fast string search. Prompts the user for a string and then displays a new buffer which contains all records whose any field contains this string.

In both of these commands, if the recfile has a record descriptor, it is copied as well.

5.2 Selection mode

In both navigation mode and the assorted edition modes, rec-cmd-goto-next-rec and rec-cmd-goto-prev-rec navigate to the next and previous record, respectively. On large files, finding a particular record using these commands only can be quite tedious. To this end, using selections, it is possible to jump to or restrict navigation to records matching selections.

Restricting the navigation commands, typically bound to n and C-c C-n (mutatis mutandis for backward movement), to records that match the current selection, can be done using what is called record selection mode.

Record selection mode is a buffer-local minor mode in which the commands for next and previous records as described above do not behave the usual way: they navigate records that only match the currently active selection.

When selection mode is entered, any associated summary buffer automatically updates its contents to match that of the current selection. Once selection mode is turned off, the summary buffer reverts back to the previous selection.

Record selection mode can be entered by the following commands:

/ s
M-x rec-cmd-navigate-current-type-by-sex

Prompts the user for a selection expression, runs a query on the current buffer and the current type, and then enters record selection mode.

/ q
M-x rec-cmd-navigate-current-type-by-fast-string

Same as above, but with a fast string search.

In record selection mode, besides the standard navigation bindings, the following additional bindings are defined:

/ /
M-x rec-cmd-exit-selection

Exits record selection mode and restores the normal navigation keybindings.

6 Cross referencing

Besides navigation, edition and selection, rec-mode provides another powerful feature for recfiles, known as cross referencing. In short, cross referencing is an Emacs term for listing various kinds of search results for certain keywords an identifiers. For records, this could mean the following:

Finding the referred record of a foreign key. If a field is a key referring to another record, you can jump to that record.

1. Finding the referents of a record if it is a foreign key.

2. Finding records that match a certain selection expression or fast string search.

Emacs provides a unified interface for this called Xref, see See Section “Xref” in GNU Emacs Manual. rec-mode implements a custom xref backend for the purpose of easier navigating between records.

Regardless of functionality, if there are more than one search result, Xref searches present their results in a special XREF buffer. It is an ordered list of results, similar to compilation mode (see Section “Compilation Mode” in GNU Emacs Manual). Xref makes it easy to navigate between search results, as it binds the next-error functionality, letting you navigate the results even if you are outside the rec mode buffer.

6.1 Cross-referencing records using foreign keys

When a record field is a reference to another record, it is possible to use navigate from the referring field to that record. Consider the following record:

%rec: Book
%type: Published int
%type: Author rec Author
Title: Cat's Cradle
Published: 1963
Author: KurtVonnegut
Title: Breakfast of Champions
Published: 1973
Author: KurtVonnegut
Title: The Left Hand of Darkness
Published: 1969
Author: UrsulaKLeGuin
%rec: Author
%key: Id
Id: UrsulaKLeGuin
Name: Ursula K. Le Guin
Id: KurtVonnegut
Name: Kurt Vonnegut, Jr.

6.1.1 Navigating to referred records

Navigating to Author field of the Book records and hitting M-. (xref-goto-definition) will jump the editor to the corresponding Author record. The requirement is that the current field is a valid foreign key, i.e. the field is a record reference and the referred record exists in the current buffer.

6.1.2 Finding record references

On the other hand, in the above example, navigating anywhere on an Author record, it is possible to list all the records referring that record using M-? (xref-find-references). This works on any field of the record, since we foreign keys to the record work on the record level.

6.1.3 Commands for cross-referencing

M-.
M-x xref-goto-definition

Jump to the record referenced by this field, if possible. Use M-. and C-M-, to jump between the results.

M-?
M-x xref-find-references

List all the records that reference this record. The record referenced must have a %key in its descriptor for this to work, and to get any results the referring records must have properly used foreign keys.

For more keybindings, See Section “Looking Up Identifiers” in Gnu Emacs Manual. For more info about foreign keys in recutils, See Section “Foreign Keys” in Recutils Manual.

6.2 Cross referencing by selection

Cross referencing by selection lets you perform a query and then list the results using the Xref buffer. Rec mode provides two such commands:

X s
M-x rec-cmd-xref-sex

Query the current recfile using a selection expression and the current type and then list the results using Xref. Each results shows a small snippet of the matched record making it easier to identify which records matched the search.

X q
M-x rec-cmd-xref-fast-string

Like above, but using a fast string. Instead of previewing the first three fields of a record, it shows each field where there was a match.

7 Configuration

rec-mode can be configured using the customization facilities of Emacs.

7.1 Finding the recutils

rec-mode makes use of the several utilities which are part of the recutils. The following variables tell the mode where to find the utilities. The default values of these variables must work if the recutils are installed system-wide in the system.

[Variable] rec-recsel

Name of the recsel utility from the GNU recutil.

[Variable] rec-recinf

Name of the recinf utility from the GNU recutils.

[Variable] rec-recfix

Name of the recfix utility from the GNU recutils.

7.2 Records appearance

The appearance of the records in navigation mode can be customised by tweaking the value of the following variables.

[Variable] rec-max-lines-in-fields

Values in fiels having more than the specified number of lines will be hidden by default in navigation mode. When hidden, an ellipsis is shown instead of the value of the field. Default is 15.

Command and Function Index

R

rec-cmd-compile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 rec-cmd-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 rec-cmd-edit-field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 rec-cmd-exit-selection . . . . . . . . . . . . . . . . . . . . . . . . . . 8 rec-cmd-goto-next-field. . . . . . . . . . . . . . . . . . . . . . . . . 2 rec-cmd-goto-next-rec . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

rec-cmd-goto-previous-rec . . . . . . . . . . . . . . . . . . . . . . 2

rec-cmd-jump-back. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 rec-cmd-jump-to-fast . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 rec-cmd-jump-to-sex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

rec-cmd-navigate-current-type-

by-fast-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

rec-cmd-navigate-current-type-by-sex. . . . . . . . . . 8
rec-cmd-new-buffer-from-fast-string . . . . . . . . . . . 7
rec-cmd-new-buffer-from-sex . . . . . . . . . . . . . . . . . . . . 7

rec-cmd-show-descriptor. . . . . . . . . . . . . . . . . . . . . . . . . 2 rec-cmd-show-info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 rec-cmd-show-summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 rec-cmd-show-type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 rec-cmd-statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

                                                  rec-cmd-toggle-field-visibility . . . . . . . . . . . . . . . 3
                                                  rec-cmd-trim-field-value . . . . . . . . . . . . . . . . . . . . . . . 3
                                                  rec-cmd-xref-fast-string . . . . . . . . . . . . . . . . . . . . . . 10

rec-cmd-xref-sex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 rec-edit-buffer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 rec-edit-record . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 rec-edit-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 rec-find-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 rec-finish-editing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 rec-summary-cmd-next . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

X

xref-find-references . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 xref-goto-definition . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

Key (Character) Index

#

# . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

%

% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

/

/ / . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 / q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 / s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

B

b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

C

c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 C-c C-c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 C-c C-t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

D

d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

E

e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

H

h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

I

I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

M

m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 M-. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 M-? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

N

n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 6

P

p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

R

R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

S

s q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 s s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 SPC. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

T

t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 TAB. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

X

x q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 X q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 x s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 X s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

Variable and Concept Index

A

activation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 author. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

B

bug reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

C

counting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 cross referencing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 cross referencing, foreign keys . . . . . . . . . . . . . . . . . . . . . . 9 cross referencing, selection . . . . . . . . . . . . . . . . . . . . . . . . 10 cross-referencing, backreferences . . . . . . . . . . . . . . . . . . 10 cross-referencing, goto-definition . . . . . . . . . . . . . . . . . . . 9 cross-referencing, keybindings . . . . . . . . . . . . . . . . . . . . . 10

E

editing fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 edition modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 edition, fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

F

feedback. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 fixing records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 folding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

J

jumping, to records . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

M

maintainer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 modes, edition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 modes, navigation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 modes, selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 modes, summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 motion, between records . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

N

navigation mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

R

rec-max-lines-in-fields . . . . . . . . . . . . . . . . . . . . . . . 11 rec-recfix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 rec-recinf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 rec-recsel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 record navigation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

S

selection mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 selection, buffer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

X

xref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008, 2020 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License.

A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other copy-

right notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item stating

at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their

titles. Section numbers or the equivalent are not considered part of the section titles. M. Delete any section Entitled “Endorsements”. Such a section may not be included in

the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in title

with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See http://

www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.